A block of mass 20kg slides downward a plane inclined at 60° with the horizontal. The coefficient friction between the plane and the block is 0.4. Find the acceleration of the block


1
Expert's answer
2021-04-15T07:17:23-0400



The forces acting on the block are shown on the picture. Along the axis, perpendicular to the surface of the plane, "N = m g \cos \theta".

According to 2nd Newton's law, the acceleration is "m a = m g \sin \theta - f", where "f" is the force of friction, "f = \mu N", which is equal to "f = \mu m g \cos \theta", using equation above.

Hence,

"m a = m g \sin \theta - \mu m g \cos \theta = m g(\sin \theta - \mu \cos \theta)", from where the acceleration is "a = g (\sin \theta - \mu \cos \theta) \approx 6.53 \frac{m}{s^2}".